One year ago, Gill was five times as old as her horse. In one year’s time, the sum of their ages will be 22 . How old was Gill three years ago?
Found 2 solutions by stanbon, macston:
Answer by stanbon(75887)

One year ago, Gill was five times as old as her horse. In one year’s time, the sum of their ages will be 22 . How old was Gill three years ago?
------------------
Equations:
g-1 = 5(h-1)
g+1 + h+1 = 22
---------------------
Rearrange::
g - 5h = -4
g + h = 20
----------------
Subtract and solve for "h":
6h = 24
h = 4 yrs (horse's age now)
----
Solve for "g"::
g = 20-h = 16 yrs (Gill's age now)
--------------
Cheers,
Stan H.
------------

Answer by macston(5194)

G=Gill's age; H=horse's age
G+1yr+H+1yr=22yrs
G+H+2yrs=22yrs
G+H=20yrs
G=20yrs-H
G-1yr=5(H-1yr) Substitute for G
(20yrs-H)-1yr=5H-5yrs
19yrs-H=5H-5yrs Add 5yrs+H to each side
24yrs=6H Divide each side by 6.
4yrs=H The horse is 4 years old.
G=20yrs-H=20yrs-4yrs=16yrs Gill is 16 years old
G-3yrs=16yrs-3yrs=13yrs ANSWER: Three years ago, Gill was thirteen years old.